### v1

# Fixed Points and Cycles of Parking Functions

Abstract.
A parking function of length $n$ is a sequence $\pi=(\pi_{1},\dots,\pi_{n})$ of positive integers such that if $\lambda_{1}\leq\cdots\leq\lambda_{n}$ is the increasing rearrangement of $\pi_{1},\dots,\pi_{n}$, then $\lambda_{i}\leq i$ for $1\leq i\leq n$. In this paper we obtain some exact results on the number of fixed points and cycles of parking functions. Our proofs will be based on generalizations of Pollak’s argument. Extensions of our techniques are discussed.
Key words and phrases:
Parking function; Pollak’s circle argument
2010 Mathematics Subject Classification:
05A15; 05A19, 60C05
M. Yin was supported by the University of Denver’s Professional Research Opportunities for Faculty Fund 80369-145601.
1. Introduction
Parking functions were introduced by Konheim and Weiss [2] in the study of the linear probes of random hashing functions. In the classical parking function scenario, we have $n$ parking spaces on a one-way street, labelled $1,2,\dots,n$ in consecutive order as we drive down the street. There are $n$ cars $C_{1},\dots,C_{n}$. Each car $C_{i}$ has a preferred space $1\leq\pi_{i}\leq n$. The cars drive down the street one at a time in the order $C_{1},\dots,C_{n}$. The car $C_{i}$ drives immediately to space $i$ and then parks in the first available space. Thus if $i$ is empty, then $C_{i}$ parks there; otherwise $C_{i}$ next goes to space $i+1$, etc. If all cars are able to park, then the sequence $\pi=(\pi_{1},\dots,\pi_{n})$ is called a parking function of length $n$. It is well-known and easy to see that if $\lambda_{1}\leq\cdots\leq\lambda_{n}$ is the (weakly) increasing rearrangement of $\pi_{1},\dots,\pi_{n}$, then $\pi$ is a parking function if and only if $\lambda_{i}\leq i$ for $1\leq i\leq n$. In particular, any permutation of a parking function is a parking function. Write $\operatorname{PF}_{n}$ for the set of parking functions of length $n$.
The first significant result on parking functions, due to Pyke [4] in another context and then to Konheim and Weiss [2], is that the number of parking functions of length $n$ is equal to $(n+1)^{n-1}$. A famous combinatorial proof was given by Pollak (unpublished but recounted in [1] and [5]). It boils down to the following easily verified statement: Let $G$ denote the group of all $n$-tuples $(a_{1},\dots,a_{n})\in[n+1]^{n}$ with componentwise addition modulo $n+1$. Let $H$ be the subgroup generated by $(1,1,\dots,1)$. Then every coset of $H$ contains exactly one parking function.
We will be interested in deriving some exact formulas for the number of fixed points and cycles of parking functions in this paper. For a parking function $\pi=(\pi_{1},\dots,\pi_{n})$, we say $i$ is a fixed point of $\pi$ if $\pi_{i}=i$. More generally, for $m\geq 1$, we say $(i_{1},\dots,i_{m})$ where the $i_{j}$’s are all distinct is an $m$-cycle of $\pi$ if $\pi_{i_{1}}=i_{2},\pi_{i_{2}}=i_{3},\dots,\pi_{i_{m-1}}=i_{m},\pi_{i_{m}}=i_{1}$, so fixed points are simply $1$-cycles. Our proofs will be based on generalizations of Pollak’s argument. In Section 2 we study classical parking functions, in Section 3 we study prime parking functions, and in Section 4 we discuss extensions of our techniques to $(r,k)$-parking functions.
2. Classical parking functions
Our first result identifies the number of parking functions with a prescribed number of fixed points.
Theorem 2.1.
Let $0\leq k\leq n$. The number of parking functions of length $n$ with $k$ fixed points is given by
$\frac{1}{(n+1)^{2}}\binom{n+1}{k}\left(n^{n-k+1}+(-1)^{n+k}\right).$
Proof.
For $1\leq i\leq n$, let $A_{i}=|\{\pi\in\operatorname{PF}_{n}:\pi_{i}=i\}|$. By De Morgan’s law and the inclusion-exclusion principle, the number of parking functions of length $n$ with $k$ fixed points is
$\hskip 56.9055pt\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\left|\left(\bigcap_{j=1}^{k}A_{i_{j}}\right)\bigcap_{i\neq i_{j}\,\forall 1\leq j\leq k}A_{i}^{c}\right|$
$=$ $\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\sum_{\ell=0}^{n-k}(-1)^{\ell}\sum_{\begin{subarray}{c}1\leq s_{1}<\cdots<s_{\ell}\leq n\\ s_{t}\neq i_{j}\,\forall 1\leq t\leq\ell\,\forall 1\leq j\leq k\end{subarray}}\left|\left(\bigcap_{j=1}^{k}A_{i_{j}}\right)\bigcap\left(\bigcap_{t=1}^{\ell}A_{s_{t}}\right)\right|$
$=$ $\sum_{\ell=0}^{n-k}(-1)^{\ell}\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\sum_{\begin{subarray}{c}1\leq s_{1}<\cdots<s_{\ell}\leq n\\ s_{t}\neq i_{j}\,\forall 1\leq t\leq\ell\,\forall 1\leq j\leq k\end{subarray}}\left|\left(\bigcap_{j=1}^{k}A_{i_{j}}\right)\bigcap\left(\bigcap_{t=1}^{\ell}A_{s_{t}}\right)\right|$
$=$ $\sum_{\ell=0}^{n-k}(-1)^{\ell}\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\sum_{\begin{subarray}{c}1\leq s_{1}<\cdots<s_{\ell}\leq n\\ s_{t}\neq i_{j}\,\forall 1\leq t\leq\ell\,\forall 1\leq j\leq k\end{subarray}}\left|\left\{\pi\in\operatorname{PF}_{n}:\begin{subarray}{c}\pi_{1}<\cdots<\pi_{k+\ell}\\ \left\{\pi_{1},\ldots,\pi_{k+\ell}\right\}=\left\{i_{1},\ldots,i_{k},s_{1},\ldots,s_{\ell}\right\}\end{subarray}\right\}\right|,$
where the last equality follows by symmetry of parking coordinates, and by
$\left\{\pi_{1},\ldots,\pi_{k+\ell}\right\}=\left\{i_{1},\ldots,i_{k},s_{1},\ldots,s_{\ell}\right\},$
we mean equality as a set.
To obtain a simplified expression for this last equation, we use an extension of Pollak’s circle argument. Add an additional parking spot $n+1$, and arrange the spots in a circle. Allow $n+1$ also as a preferred spot. We first select $k+\ell$ spots for the first $k+\ell$ cars. Note that it matters whether these spots correspond to $i_{j}$ for $1\leq j\leq k$ or to $s_{t}$ for $1\leq t\leq\ell$, and altogether this can be done in $\binom{n+1}{k,\ell,n-k-\ell+1}$ ways. Then for the remaining $n-k-\ell$ cars, there are $(n+1)^{n-k-\ell}$ possible preference sequences. Out of the $n+1$ rotations for any preference sequence, only one rotation becomes a valid parking function. Standard circular symmetry argument yields
$\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\sum_{\begin{subarray}{c}1\leq s_{1}<\cdots<s_{\ell}\leq n\\ s_{t}\neq i_{j}\,\forall 1\leq t\leq\ell\,\forall 1\leq j\leq k\end{subarray}}\left|\left\{\pi\in\operatorname{PF}_{n}:\begin{subarray}{c}\pi_{1}<\cdots<\pi_{k+\ell}\\ \left\{\pi_{1},\ldots,\pi_{k+\ell}\right\}=\left\{i_{1},\ldots,i_{k},s_{1},\ldots,s_{\ell}\right\}\end{subarray}\right\}\right|$
$\hskip 56.9055pt=\binom{n+1}{k,\ell,n-k-\ell+1}(n+1)^{n-k-\ell-1}.$
So what we need to compute is
$\sum_{\ell=0}^{n-k}(-1)^{\ell}\binom{n+1}{k,\ell,n-k-\ell+1}(n+1)^{n-k-\ell-1}=\binom{n+1}{k}\frac{((n+1)-1)^{n-k+1}+(-1)^{n+k}}{(n+1)^{2}}.$
Extending the notion of fixed points to cycles, we have a more general result.
Theorem 2.2.
Let $m\geq 1$ and $k\geq 0$ with $km\leq n$. The number of parking functions of length $n$ with $k$ $m$-cycles is given by
$\sum_{\ell:(k+\ell)m\leq n}(-1)^{\ell}\frac{\left((m-1)!\right)^{k+\ell}}{k!\ell!}\binom{n+1}{\underbracket{m,\cdots,m}_{k+\ell\hskip 2.84544pt\text{$m$'s}},n-(k+\ell)m+1}(n+1)^{n-(k+\ell)m-1}.$
Proof.
The proof follows similarly as in Theorem 2.1. We will not include all the technical details, but instead walk through the key ideas.
For distinct $1\leq i_{1},\dots,i_{m}\leq n$, let
$A_{(i_{1},\dots,i_{m})}=|\{\pi\in\operatorname{PF}_{n}:\pi_{i_{1}}=i_{2},\pi_{i_{2}}=i_{3},\dots,\pi_{i_{m-1}}=i_{m},\pi_{i_{m}}=i_{1}\}|.$
We still apply De Morgan’s law, the inclusion-exclusion principle, and an extension of Pollak’s circle argument. The major difference is that instead of selecting $k+\ell$ spots on a circle with $n+1$ spots, we select $k+\ell$ $m$-tuples on the circle. We also include an additional scalar factor $(m-1)!$ for each such tuple, which accounts for the number of ways of arranging an $m$-cycle with distinct entries. ∎
As is standard, we denote the probability distribution and expectation with respect to the uniform measure on the set of classical parking functions of length $n$ by $\mathbb{P}_{n}$ and $\mathbb{E}_{n}$, respectively. The following theorem provides an exact result on the expected number of $m$-cycles in a classical parking function drawn uniformly at random.
Theorem 2.3.
Take $1\leq m\leq n$. Let $\pi\in\operatorname{PF}_{n}$ be a parking function chosen uniformly at random and $C_{m}(\pi)$ be the number of $m$-cycles of $\pi$. The expected number of $m$-cycles is given by
$\mathbb{E}_{n}(C_{m}(\pi))=(m-1)!\binom{n+1}{m}/(n+1)^{m}.$
Remark 2.4.
From Theorem 2.3, we may readily derive that for fixed $m$, as $n$ gets large, the expected number of $m$-cycles in a random parking function is asymptotically $1/m$. This asymptotic result was obtained earlier by Paguyo [3] using other means.
Proof.
For distinct $1\leq i_{1},\dots,i_{m}\leq n$, let
$A_{(i_{1},\dots,i_{m})}=|\{\pi\in\operatorname{PF}_{n}:\pi_{i_{1}}=i_{2},\pi_{i_{2}}=i_{3},\dots,\pi_{i_{m-1}}=i_{m},\pi_{i_{m}}=i_{1}\}|.$
By linearity of expectation and symmetry of parking coordinates,
$\mathbb{E}_{n}(C_{m}(\pi))$ $=\sum_{1\leq i_{1}<\cdots<i_{m}\leq n}(m-1)!\,\mathbb{P}_{n}(A_{(i_{1},\dots,i_{m})})$
$=\sum_{1\leq i_{1}<\cdots<i_{m}\leq n}(m-1)!\,\mathbb{P}_{n}(\pi_{1}=i_{1},\dots,\pi_{m}=i_{m})$
$=\frac{(m-1)!}{(n+1)^{n-1}}\,\left|\left\{\pi\in\operatorname{PF}_{n}:\pi_{1}<\pi_{2}<\cdots<\pi_{m}\right\}\right|,$
where the additional scalar factor $(m-1)!$ in the first equation accounts for the number of ways of arranging an $m$-cycle with distinct entries $i_{1},\dots,i_{m}$.
To obtain a simplified expression for this last equation, we use an extension of Pollak’s circle argument as in the proof of Theorem 2.1. Add an additional parking spot $n+1$, and arrange the spots in a circle. Allow $n+1$ also as a preferred spot. We first select $m$ spots for the first $m$ cars, which can be done in $\binom{n+1}{m}$ ways. Then for the remaining $n-m$ cars, there are $(n+1)^{n-m}$ possible preference sequences. Out of the $n+1$ rotations for any preference sequence, only one rotation becomes a valid parking function. Standard circular symmetry argument yields
|$\left|\left\{\pi\in\operatorname{PF}_{n}:\pi_{1}<\pi_{2}<\cdots<\pi_{m}\right\}\right|=\binom{n+1}{m}(n+1)^{n-m-1}.$
Our conclusion follows. ∎
3. Prime parking functions
A classical parking function $\pi=(\pi_{1},\dots,\pi_{n})$ is said to be prime if for all $11\leq j\leq n-$ 1, at least $j+1$ cars want to park in the first $j$ places. (Equivalently, if we remove some term of $\pi$ equal to 1, then we still have a parking function.) Denote the set of prime parking functions of length $n$ by $\operatorname{PPF}_{n}$. Note that for prime parking functions it is impossible to have $n$ fixed points or to have an $n$-cycle.
As with classical parking functions, we can also study prime parking functions via circular rotation. The modified circular symmetry argument was first due to Kalikow [7, pp. 141-142], who provided the following observation: Let $G$ denote the group of all $n$-tuples $(a_{1},\dots,a_{n})\in[n-1]^{n}$ with componentwise addition modulo $n-1$. Let $H$ be the subgroup generated by $(1,1,\dots,1)$. Then every coset of $H$ contains exactly one prime parking function. The results we present in this section are thus largely parallel to the results for classical parking functions in our previous section.
Theorem 3.1.
Let $10\leq k\leq n-$ 1. The number of prime parking functions of length $n$ with $k$ fixed points is given by
$\binom{n-1}{k}(n-2)^{n-k-1}.$
Proof.
We proceed as in the proof of Theorem 2.1. The only difference is that the circular symmetry argument is now for a circle with $n-1$ spots. So we have
$\sum_{\ell=0}^{n-k-1}(-1)^{\ell}\binom{n-1}{k,\ell,n-k-\ell-1}(n-1)^{n-k-\ell-1}=\binom{n-1}{k}(n-2)^{n-k-1}.$
As in the case of classical parking functions, a more general theorem exists for prime parking functions when we extend the notion of fixed points to cycles.
Theorem 3.2.
Let $m\geq 1$ and $k\geq 0$ with $km\leq n-1$. The number of prime parking functions of length $n$ with $k$ $m$-cycles is given by
$\sum_{\ell:(k+\ell)m\leq n-1}(-1)^{\ell}\frac{\left((m-1)!\right)^{k+\ell}}{k!\ell!}\binom{n-1}{\underbracket{m,\cdots,m}_{k+\ell\hskip 2.84544pt\text{$m$'s}},n-(k+\ell)m-1}(n-1)^{n-(k+\ell)m-1}.$
Proof.
We proceed as in the proof of Theorem 2.2 and note that the circular symmetry argument is now for a circle with $n-1$ spots. ∎
As is standard, we denote the probability distribution and expectation with respect to the uniform measure on the set of prime parking functions of length $n$ by $\mathbb{P}_{n}$ and $\mathbb{E}_{n}$, respectively. The following theorem provides an exact result on the expected number of $m$-cycles in a prime parking function drawn uniformly at random.
Theorem 3.3.
Take $11\leq m\leq n-$ 1. Let $\pi\in\operatorname{PPF}_{n}$ be a prime parking function chosen uniformly at random and $C_{m}(\pi)$ be the number of $m$-cycles of $\pi$. The expected number of $m$-cycles is given by
$\mathbb{E}_{n}(C_{m}(\pi))=(m-1)!\binom{n-1}{m}/(n-1)^{m}.$
Remark 3.4.
From Theorem 3.3, we may readily derive that for fixed $m$, as $n$ gets large, the expected number of $m$-cycles in a random prime parking function is asymptotically $1/m$.
Proof.
We proceed as in the proof of Theorem 2.3. The only difference is that the circular symmetry argument is now for a circle with $n-1$ spots. So we have
$\mathbb{E}_{n}(C_{m}(\pi))=\frac{(m-1)!}{(n-1)^{n-1}}\binom{n-1}{m}(n-1)^{n-m-1}=(m-1)!\binom{n-1}{m}/(n-1)^{m}.$
4. Further discussions
Fix some positive integers $k$ and $r$. Define an $(r,k)$-parking function of length $n$ [8] to be a sequence $(\pi_{1},\dots,\pi_{n})$ of positive integers such that if $\lambda_{1}\leq\cdots\leq\lambda_{n}$ is the (weakly) increasing rearrangement of $\pi_{1},\dots,\pi_{n}$, then $\lambda_{i}\leq k+(i-1)r$ for $1\leq i\leq n$. Denote the set of $(r,k)$-parking functions of length $n$ by $\operatorname{PF}_{n}(r,k)$. There is a similar interpretation for such $(r,k)$-parking functions in terms of the classical parking function scenario: One wishes to park $n$ cars on a street with $k+(n-1)r$ spots, but only $n$ spots are still empty, which are at positions no later than $k,\dots,k+(n-1)r$.
Pollak’s original circle argument [1] may be further extended to $(r,k)$-parking functions. Let $G$ denote the group of all $n$-tuples $(a_{1},\dots,a_{n})\in[k+nr]^{n}$ with componentwise addition modulo $k+nr$. Let $H$ be the subgroup generated by $(1,1,\dots,1)$. Then every coset of $H$ contains exactly $k$ $(r,k)$-parking functions. See [6] for more details.
The following proposition is then immediate.
Proposition 4.1.
Let $0\leq s\leq n$. We have
|$|\{\pi\in\operatorname{PF}_{n}(r,k):\pi_{1}<\pi_{2}<\dots<\pi_{s}\}|=k\binom{k+nr}{s}(k+nr)^{n-s-1}.$
Proof.
We apply the generalized circle argument as described above. Arrange $k+nr$ spots in a circle. We first select $s$ spots for the first $s$ cars, which can be done in $\binom{k+nr}{s}$ ways. Then for the remaining $n-s$ cars, there are $(k+nr)^{n-s}$ possible preference sequences. Out of the $k+nr$ rotations for any preference sequence, exactly $k$ rotations become valid $(r,k)$-parking functions. The rest is standard circular symmetry argument. ∎
acknowledgements
We thank Richard Stanley for helpful discussions. Some of the results in this paper were independently conjectured by Ari Cruz.